# A Survey on Class-Agnostic Counting: Advancements from Reference-Based to Open-World Text-Guided Approaches

### v3

@@ Abstract
Visual object counting has recently shifted towards class-agnostic counting (CAC), which addresses the challenge of counting objects across arbitrary categories – a crucial capability for flexible and generalizable counting systems. Unlike humans, who effortlessly identify and count objects from diverse categories without prior knowledge, most existing counting methods are restricted to enumerating instances of known classes, requiring extensive labeled datasets for training and struggling in open-vocabulary settings. In contrast, CAC aims to count objects belonging to classes never seen during training, operating in a few-shot setting. In this paper, we present the first comprehensive review of CAC methodologies. We propose a taxonomy to categorize CAC approaches into three paradigms based on how target object classes can be specified: reference-based, reference-less, and open-world text-guided. Reference-based approaches achieve state-of-the-art performance by relying on exemplar-guided mechanisms. Reference-less methods eliminate exemplar dependency by leveraging inherent image patterns. Finally, open-world text-guided methods use vision-language models, enabling object class descriptions via textual prompts, offering a flexible and promising solution. Based on this taxonomy, we provide an overview of the architectures of 29 CAC approaches and report their results on gold-standard benchmarks. We compare their performance and discuss their strengths and limitations. Specifically, we present results on the FSC-147 dataset, setting a leaderboard using gold-standard metrics, and on the CARPK dataset to assess generalization capabilities. Finally, we offer a critical discussion of persistent challenges, such as annotation dependency and generalization, alongside future directions. We believe this survey will serve as a valuable resource for researchers, highlighting the progressive developments in CAC and offering insights for future advancements in the field.
Index Terms:
Object Counting, Class-agnostic Counting, Few-shot Counting, Prompt-based Counting, Deep Learning, Survey
publicationid: pubid: Preprint. Under review.
@@ I Introduction
The goal of object counting is to automatically estimate the number of object instances in still images or video frames [1]. This challenging task has emerged as a prominent research focus within the realm of computer vision, owing to its vast array of practical applications and interdisciplinary significance. For instance, crowd counting [2, 3, 4, 5, 6, 7, 8, 9, 10] is crucial in event management and analysis, holding paramount importance from social [11], psychological [12], political [13], and security [14] perspectives. In smart agriculture, object counting is vital for livestock monitoring [15, 16], pest abundance estimation [17], and for assessing seedling emergence rates by counting plants and trees [18, 19, 20]. Counting techniques are also valuable in traffic monitoring [21, 22, 23, 24, 25], inventory management [26], biomedical applications [27, 28, 29, 30], and ecological surveys [31, 32].
Figure 1: Comparison between class-specific and class-agnostic counting. On the left, we report two examples of class-specific counters, which need individually trained networks on extensive labeled datasets for each object type. On the right, we show the new class-agnostic counting setting, where the counter estimates the number of objects belonging to arbitrary classes. Note that, in the latter scenario, training and test object classes are different.
However, while humans can effortlessly count objects from various visual categories, current state-of-the-art AI-based methods generally focus on enumerating the instances of a single or small set of known classes. Indeed, humans possess an intuitive ability to discern what merits counting when confronted with unfamiliar objects without requiring a reference example or prior comprehension of the object type. In essence, this inherent counting ability comprises two primary facets: recognizing what is noteworthy for counting and detecting repetitions of countable elements. In contrast, even the very recent general-purpose AI agents and Multi-modal Large Language Models (MLLMs) such as LLaVA [33], MiniGPT4 [34], and mPLUG-owl [35], perform poorly on counting objects belonging to open-vocabulary classes [36], and most state-of-the-art counting approaches are class-specific, requiring individually trained networks and, consequently, large labeled datasets for each object type [37] (see left part of Fig. 1).
Figure 2: Overview of class-agnostic counting paradigms. We propose a taxonomy to classify existing CAC methodologies: a) Reference-based approaches rely on annotated bounding box exemplars, which serve as visual prototypes for the object classes to be counted; b) Reference-less techniques relax these requirements, enabling models to automatically identify the dominant class(es) to be counted; and c) Open-world Text-guided methodologies allow the use of textual descriptions to specify the object classes to be considered.
@@
Very recently, research about object counting has shifted towards a new trend that aims to consider arbitrary object categories and attempts to alleviate the problem of the annotation burden (see right part of Fig. 1). The first tentative of creating a model able to count any class, i.e., to perform Class-Agnostic Counting (CAC), is represented by the work of Lu and Zisserman that proposed a Generic Matching Network (GMN) [38]. To achieve this goal, they formulated counting as a matching problem, leveraging the inherent self-similarity property of images, as objects to be counted within an image are typically repeated. GMN was pre-trained with tracking video data, and it exploits an adaptation module to adapt it to domains of interest. However, the latter adaptation module needs dozens to hundreds of examples and, without adaptation, does not perform well in novel classes.
The work that has undergone a detour from standard class-specific counting is [37]. Here, the authors treated counting as a few-shot regression task. Specifically, in this new setting, the inputs are an image and some so-called exemplars; the latter are defined as bounding boxes around the objects of interest from the same input image and are used by the network as object prototypes. At inference time, similar to the few-shot classification task convention [39, 40], the classes of the objects to be counted are entirely different from the ones seen during training, i.e., the provided exemplars belong to arbitrarily unseen categories. This makes few-shot counting very different from the typical counting task, where the training and test classes are the same. It is worth noting that this paradigm does not explicitly learn from limited data – e.g., using (unsupervised) domain adaptation [41, 42] or uncertainty quantification [43]. Instead, this strategy involves using some data in a fully supervised fashion to learn a class-agnostic counting model that can later be applied to unseen object categories. In this way, it is a data-efficient methodology with respect to the unseen object categories.
@@
For the first time, this survey explores and analyzes the leading works existing in the literature tackling class-agnostic counting. Although this task is very recent, it has been addressed by several approaches published in top-tier venues, leading to increasingly better results. Furthermore, its formulation has slightly changed and evolved compared to its original version, progressively reducing the human intervention in the overall process, i.e., the need for annotated exemplars, either at training or inference time. In this context, our contributions are fourfold: (i) we propose a taxonomy to categorize CAC methods, bringing clarity to the literature; (ii) we provide an overview of the architectures of 29 CAC approaches; (iii) we present and compare the performance of the examined techniques on gold-standard benchmarks; (iv) we critically discuss the results from multiple perspectives.
@@
Specifically, the proposed taxonomy organizes the surveyed CAC methods into three categories, also illustrated in Fig. 2. The first category follows the previously mentioned reference-based or few-shot paradigm [37], where annotated bounding box exemplars are required during both training and inference. Subsequently, CAC evolved into a formulation that shapes our second category, where manually specifying object categories is no longer necessary; the algorithm automatically identifies the dominant class(es) to be counted. This approach is known in the literature as reference-less or exemplar-free. Finally, the most recent paradigm, referred to as open-world text-guided, prompt-based, or zero-shot, allows the system to take natural language descriptions of object classes as input, going hand in hand with the latest vision-language foundation models, and constitutes our third category. Based on this taxonomy, we present an overview of the architectures of the examined methods and show their performance on established gold-standard benchmarks from the literature. Specifically, we consider the FSC-147 dataset [37], which includes over 6,000 images across 147 object categories, and the CARPK dataset [44], explicitly designed for vehicle counting and used to assess cross-dataset generalization. Lastly, we critically discuss the results, highlighting the strengths and limitations of the considered methodologies, the persistent challenges in the field, and potential directions for future research.
@@
Figure 3: Overview of selected CAC methodologies over time. We highlight key milestone methods along a temporal line. Colors represent their category based on our taxonomy: red for reference-based, green for reference-less, and blue for open-world text-guided CAC approaches.
@@ II Existing Surveys, Scope, and Search Strategy
While there are several survey papers reviewing counting methods present in the computer vision literature, most of them concentrate on specific object categories, with a particular emphasis on crowd counting. The most cited survey is [45]. In this work, Sindagi et al. revised the papers up to 2017 in detail, focusing on crowd counting and considering both traditional methodologies and techniques based on convolutional neural networks. Furthermore, they described the most used crowd counting datasets, presenting and discussing the results obtained against these benchmarks. More recently, Li et al. published a similar, more updated survey in [46], retracing the last 20 years of object counting, mainly focusing on crowd counting with density estimation using convolutional neural networks. A plethora of other surveys or short surveys have been published over the last few years: a non-exhaustive list includes [47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. Even if the authors of these works look from different angles the topic, considering different taxonomies and perspectives, they always focus on single-class object counting.
Conversely, in this survey, we review the works in the literature concerning class-agnostic object counting, which refers to a new, recent, and different formulation of the classic object counting task. Here, the goal is to count objects belonging to arbitrary categories never seen during the training stage, and this new paradigm brings along completely different solutions, including different methodologies, architectures, and benchmarks. To the best of our knowledge, the only other existing survey in the literature that addresses this topic is [57]. However, in [57], CAC is treated marginally, and the authors did not cover all the existing literature but just a few works, also because the task was just born at the time of writing.
To select papers for this survey, we began by identifying the gold-standard benchmarks for the CAC task. We then conducted a comprehensive search on Google Scholar for papers that cited these benchmarks. To ensure the quality and reliability of the sources, we included only peer-reviewed published papers in our selection. We excluded non-peer-reviewed preprints from arXiv, except for those that offered freely available implementations that could guarantee the reproducibility of their results.
@@
TABLE I: List of the considered methods. We summarize the methodologies discussed in this survey, along with some relevant associated details. The methods are categorized based on our proposed taxonomy – reference-based, reference-less, and open-world text-guided.
@@
Method Venue Year Code Point-level Supervision Test-time Adaptation
@@
III Existing Approaches
In the following, we review the considered methodologies, categorizing them into three sections based on our proposed taxonomy: reference-based, reference-less, and open-world text-guided approaches. Some techniques incorporate additional modules or slight architectural modifications from their original versions, leading to potential overlaps across these groups. However, we classify each method under its primary category while describing its extensions in other relevant sections as needed. Table I summarizes the surveyed techniques, also providing additional details such as the publication venue and year. Furthermore, Figure 3 presents an overview of selected methods that we consider key milestones, highlighting their relevance over time.
III-A Problem Formulation
State-of-the-art performance in standard class-specific object counting is achieved using supervised deep learning methods, which typically follow two main strategies: counting through detection and counting through regression. The detection-based approach requires prior identifying individual object instances using an object detector [24, 25, 16, 10]. In contrast, regression-based methods learn a mapping between image features and the object count, bypassing the need for explicit localization. This can be achieved directly or, more commonly, through estimating a density map, i.e., a real-valued function. The final estimation of the number of object instances is achieved by summing up the pixel values of the latter [1, 19, 31, 32].
Detectors, while conceptually simpler and capable of handling a wide range of object classes due to the abundance of suitable annotated datasets, struggle with occluded objects. This limitation is the primary reason density-based approaches have largely replaced detection-based techniques. Thus, also most class-agnostic counting techniques follow the density-based paradigm, and CAC existing benchmarks reflect this in terms of label availability. Precisely, the gold standard labels needed for the supervised training of density regressors are dots (points), i.e., ground truth corresponds to binary spatial images with 1’s at the objects’ center location, indicating their existence, and 0’s at other locations without the objects. Since it has been proved to be difficult to train a density regressor with a loss directly based on dot annotations, most existing works generate smoothed target density maps by superimposing Gaussian kernels $G_{\sigma}$ centered in these dot-annotated locations and using a fixed or adaptive spread parameter $\sigma$, depending on the typical object size in the dataset [86, 1].
@@ III-A Problem Formulation
Formally, we assume to have a collection of $N$ annotated images denoted as $\mathcal{X}={(I_{1},P_{1}),\dots,(I_{N},P_{N})}$, where $I_{i}$ is the $i$-th image and $P_{i}\in\mathbb{R}^{2}$ the set of 2D point coordinates roughly localizing the objects to be counted in image $I_{i}$. The goal is to learn a regression model $f_{\theta}$ producing a density map $D^{map}=f_{\theta}(I)\in\mathbb{R}^{H\times W}$. The number of objects $n$ in an image sub-region $S\subseteq I$ is estimated by integrating $D^{map}$ over $S$, i.e., summing up pixel values in the considered region, $n=\sum_{p\in S}D^{map}_{p}$. The loss function to train the regression model is usually represented by per-pixel functions, such as minimizing the mean squared error between the ground truth and the predicted density maps.
In the few-shot – or reference-based – counting setting, for each image $I_{i}$, we also assume to have a set of $K$ exemplars expressed as bounding box coordinates $B^{E}=\{b_{i}\}_{i=1:k}\in\mathbb{R}^{4}$, denoting single samples depicting the object class to be counted from the same image $I_{i}$ [37]. In the original formulation, $K$ is set to 3 [37]. Furthermore, in few-shot counting, object classes are divided into base classes $C_{b}$ and novel classes $C_{n}$, where $C_{b}$ and $C_{n}$ have no intersection. For each image $I$ from $C_{b}$, both ground truth density maps and exemplars are provided. In contrast, only exemplars are available for images belonging to $C_{n}$, and the goal is to count objects from $C_{n}$ by leveraging the generalization knowledge from $C_{b}$. Instead, in the reference-less and text-guided zero-shot formulations, exemplar bounding boxes are not used. In the reference-less case, the algorithm automatically selects the dominant class(es) to count by analyzing the repetition of object instances. Rather, in the text-guided zero-shot, for each image $I_{i}$, we also assume to have a textual description $T_{i}$ of the object class to be counted.

### v4

@@ Abstract
Visual object counting has recently shifted towards class-agnostic counting (CAC), which addresses the challenge of counting objects across arbitrary categories—a crucial capability for flexible and generalizable counting systems. Unlike humans, who effortlessly identify and count objects from diverse categories without prior knowledge, most existing counting methods are restricted to enumerating instances of known classes, requiring extensive labeled datasets for training and struggling in open-vocabulary settings. In contrast, CAC aims to count objects belonging to classes never seen during training, operating in a few-shot setting. In this paper, we present the first comprehensive review of CAC methodologies. We propose a taxonomy to categorize CAC approaches into three paradigms based on how target object classes can be specified: reference-based, reference-less, and open-world text-guided. Reference-based approaches achieve state-of-the-art performance by relying on exemplar-guided mechanisms. Reference-less methods eliminate exemplar dependency by leveraging inherent image patterns. Finally, open-world text-guided methods use vision-language models, enabling object class descriptions via textual prompts, offering a flexible and promising solution. Based on this taxonomy, we provide an overview of 30 CAC architectures and report their performance on gold-standard benchmarks, discussing key strengths and limitations. Specifically, we present results on the FSC-147 dataset, setting a leaderboard using gold-standard metrics, and on the CARPK dataset to assess generalization capabilities. Finally, we offer a critical discussion of persistent challenges, such as annotation dependency and generalization, alongside future directions. We believe this survey offers a valuable resource for researchers, capturing the evolution of CAC and providing insights to guide future developments in the field.
@@ I Introduction
The goal of object counting is to automatically estimate the number of object instances in still images or video frames [1]. This challenging task has emerged as a prominent research focus within the realm of computer vision, owing to its vast array of practical applications and interdisciplinary significance. For instance, in crowd analysis [2, 3, 4, 5, 6, 7, 8, 9, 10, 11], it supports event monitoring with implications across social [12], psychological [13], political [14], and security domains [15]. In agriculture, it enables livestock and pest monitoring [16, 17, 18], and plant emergence assessment [19, 20, 21]. Other applications include traffic analysis [22, 23, 24, 25, 26], inventory control [27], and biomedical imaging [28, 29, 30, 31]. However, unlike humans—who can naturally identify what is worth counting and recognize repetitions even among unfamiliar objects—most AI-based methods are constrained to predefined categories. These approaches typically rely on prior knowledge of the object class and require dedicated models trained on large annotated datasets [32] (see also the left part of Fig. 1). Even recent general-purpose agents and MLLMs, such as LLaVA [33], MiniGPT4 [34], and mPLUG-owl [35], struggle with counting objects from open-vocabulary categories [36].
Figure 1: Comparison between class-specific and class-agnostic counting. On the left, we show examples of class-specific counting, where each object category (e.g., cars, people, animals) requires a dedicated model trained on large, annotated datasets. These models are tailored to recognize and count only the specific class they were trained on. On the right, we present the class-agnostic counting approach, which aims to estimate the number of objects regardless of their semantic category. In this setting, the counting model is trained on a set of object classes and evaluated on entirely different, unseen classes, demonstrating its ability to generalize across categories. This paradigm significantly reduces the need for exhaustive per-class annotations and enables broader applicability in real-world scenarios.
@@
Very recently, research about object counting has shifted towards a new trend that aims to consider arbitrary object categories and attempts to alleviate the problem of the annotation burden (see right part of Fig. 1). The first attempt of creating a model able to count any class, i.e., to perform Class-Agnostic Counting (CAC), was proposed by Lu and Zisserman through the Generic Matching Network (GMN) [37], which formulates counting as a matching task by exploiting image self-similarity—since objects to be counted often repeat within the same image. GMN is pre-trained on tracking data and includes an adaptation module for domain transfer. However, this module requires dozens to hundreds of examples and performs poorly on novel classes without adaptation. The first work to significantly depart from standard class-specific counting was introduced by Ranjan et al. [32], who introduced a few-shot regression approach to counting. In this setting, the inputs are an image and a set of exemplars—bounding boxes around objects of interest from the same image—used as object prototypes. At inference time, similar to the few-shot classification paradigm [38, 39], the objects to be counted belong to unseen categories, making few-shot counting fundamentally different from traditional counting, where training and test classes match. Notably, this paradigm does not learn from limited data via domain adaptation [40, 41] or uncertainty modeling [42], but instead uses supervised data to train a class-agnostic model applicable to novel categories—making it data-efficient for unseen classes.
Figure 2: Overview of class-agnostic counting paradigms. We propose a taxonomy to classify existing CAC methodologies: a) Reference-based approaches rely on annotated bounding box exemplars, which serve as visual prototypes for the object classes to be counted; b) Reference-less techniques relax these requirements, enabling models to automatically identify the dominant class(es) to be counted; and c) Open-world Text-guided methodologies allow the use of textual descriptions to specify the object classes to be considered. The arrows indicate the conceptual evolution of CAC paradigms, highlighting the progressive reduction of human supervision and the increasing flexibility and semantic control.
@@
For the first time, this survey explores and analyzes the leading works existing in the literature tackling class-agnostic counting. Despite being a recent task, it has already attracted attention from top-tier venues, yielding progressively better results. Furthermore, its formulation has slightly changed and evolved compared to its original version, gradually reducing human intervention in the overall process, i.e., the need for annotated exemplars, either at training or inference time. In this context, our contributions are fourfold: (i) we propose a taxonomy to categorize CAC methods, bringing clarity to the literature; (ii) we provide an overview of the architectures of 30 CAC approaches; (iii) we present and compare the performance of the examined techniques on gold-standard benchmarks; (iv) we critically discuss the results from multiple perspectives.
@@
Specifically, the proposed taxonomy organizes the surveyed CAC methods into three categories, also illustrated in Fig. 2. The first category follows the previously mentioned reference-based or few-shot paradigm [32], where annotated bounding box exemplars are required during both training and inference. Subsequently, CAC evolved into a formulation that shapes our second category, where manually specifying object categories is no longer necessary; the algorithm automatically identifies the dominant class(es) to be counted. This approach is known in the literature as reference-less or exemplar-free. Finally, the most recent paradigm, referred to as open-world text-guided, prompt-based, or zero-shot, allows the system to take natural language descriptions of object classes as input, going hand in hand with the latest vision-language foundation models, and constitutes our third category. These three paradigms reflect the evolution of CAC methodologies toward reduced human supervision and increased flexibility. Reference-based methods, while effective, are not well-suited for autonomous systems, as they require users to manually provide exemplar inputs for each new category at inference time. Reference-less approaches address this limitation by removing the need for exemplars altogether, inferring the dominant object class directly from image patterns. However, they still lack semantic control and cannot incorporate user intent, making them suitable only when the target class aligns with the most visually repetitive category in the scene. Text-guided methods extend this trajectory by introducing a more natural and flexible way to specify the target category—closer to how humans describe objects—while leveraging the semantic capabilities of textual prompts. Rather than forming isolated categories, each paradigm represents a conceptual step forward, progressively addressing the limitations of the previous one and contributing to a more flexible and interaction-aware framework for CAC.
Based on this taxonomy, we present an overview of the architectures of the examined methods and show their performance on established benchmarks from the literature. In particular, we consider the FSC-147 dataset [32], introduced together with the few-shot CAC formulation and now regarded as the gold-standard benchmark for this task. FSC-147 includes over 6,000 images across 147 object categories. We also consider the CARPK dataset [43], explicitly designed for vehicle counting and used to assess cross-dataset generalization. Lastly, we critically discuss the results, highlighting the strengths and limitations of the considered methodologies, the persistent challenges in the field, and potential directions for future research.
@@ II Existing Surveys, Scope, and Search Strategy
While there are several survey papers reviewing counting methods present in the computer vision literature, most of them concentrate on specific object categories, with a particular emphasis on crowd counting. The most cited survey is [44]. In this work, Sindagi et al. revised the papers up to 2017 in detail, focusing on crowd counting and considering both traditional methodologies and techniques based on convolutional neural networks. Furthermore, they described the most used crowd counting datasets, presenting and discussing the results obtained against these benchmarks. More recently, Li et al. published a similar, more updated survey in [45], retracing the last 20 years of object counting, mainly focusing on crowd counting with density estimation using convolutional neural networks. A plethora of other surveys or short surveys have been published over the last few years: a non-exhaustive list includes [46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. Even if the authors of these works look at the topic from different angles, considering various taxonomies and perspectives, they consistently focus on single-class object counting.
In contrast, our survey is the first to provide a comprehensive and structured overview of CAC methodologies. To the best of our knowledge, the only other survey that briefly touches on CAC is [56], which covers only a few methods—likely due to the early stage of the task at the time of writing—and does not offer a systematic taxonomy or performance analysis. Conversely, our survey introduces several distinct and novel contributions: it organizes CAC methods according to an original taxonomy, systematically analyzes state-of-the-art approaches, compares their performance across benchmarks, and discusses their limitations and open challenges.
To select papers for this survey, we conducted a comprehensive search on Google Scholar for works citing FSC-147, the gold-standard benchmark introduced together with the few-shot CAC formulation by Ranjan et al. [32] (as mentioned in Sec. I). The search covered publications up to the end of 2024; the survey was completed in January 2025 before submission. Only peer-reviewed published papers were considered, excluding non-peer-reviewed preprints from arXiv unless they provided freely available implementations to ensure reproducibility.
Figure 3: Overview of selected CAC methodologies over time.We present a chronological overview of representative Class-Agnostic Counting (CAC) methods, highlighting milestone contributions from 2021 to 2024. Each method is encoded using both color and arrow style according to the taxonomy introduced in Fig. 2: red dashed arrows for reference-based approaches that rely on annotated exemplars; green dash-dot arrows for reference-less methods that infer dominant object classes without manual input; and blue solid arrows for open-world text-guided approaches that use natural language prompts to specify target classes. This timeline illustrates the evolution of CAC paradigms over time, showing the shift from exemplar-dependent models to more flexible and generalizable solutions.
III Existing Approaches
In the following, we review existing methodologies by grouping them into three categories based on our proposed taxonomy: reference-based, reference-less, and open-world text-guided approaches. While some methods include additional modules or architectural tweaks that may span multiple categories, each is classified under its primary type, with cross-category features discussed where relevant. Table I summarizes the surveyed techniques, while Fig. 3 highlights key milestones over time.
III-A Problem Formulation
State-of-the-art class-specific counting relies on supervised deep learning, typically following two strategies: detection-based and regression-based. Detection approaches identify individual instances using object detectors [25, 26, 17, 10], while regression methods estimate object counts by learning a mapping from image features to a density map [1, 20, 57, 58], whose pixel values are integrated to obtain the final count. Detectors, while conceptually simpler and capable of handling a wide range of object classes due to the abundance of suitable annotated datasets, struggle with occluded objects. This limitation is the primary reason density-based approaches have largely replaced detection-based techniques. Thus, also most class-agnostic counting techniques follow the density-based paradigm, and CAC existing benchmarks reflect this in terms of label availability. Precisely, the gold standard labels needed for the supervised training of density regressors are dots (points), i.e., ground truth corresponds to binary spatial images with 1’s at the center location of the objects, indicating their existence, and 0’s at other locations without the objects. Since directly optimizing a loss on these sparse labels is challenging, most methods generate smoothed density maps by placing Gaussian kernels $G_{\sigma}$ at each dot location, using a fixed or adaptive $\sigma$ based on object size [59, 1].
@@ III-A Problem Formulation
TABLE I: List of the considered methods. We summarize the methodologies discussed in this survey, along with some relevant associated details. The methods are categorized based on our proposed taxonomy—reference-based, reference-less, and open-world text-guided
@@
Method Venue Year Code Pretrained Model Point-level Supervision Test-time Adaptation
@@
Formally, we assume to have a collection of $N$ annotated images denoted as $\mathcal{X}={(I_{1},P_{1}),\dots,(I_{N},P_{N})}$, where $I_{i}$ is the $i$-th image and $P_{i}\in\mathbb{R}^{2}$ the set of 2D point coordinates roughly localizing the objects to be counted in image $I_{i}$. The goal is to learn a regression model $f_{\theta}$ producing a density map $D^{map}=f_{\theta}(I)\in\mathbb{R}^{H\times W}$. The number of objects $n$ in an image sub-region $S\subseteq I$ is estimated by integrating $D^{map}$ over $S$, i.e., summing up pixel values in the considered region, $n=\sum_{p\in S}D^{map}_{p}$. The regression model is typically trained by minimizing per-pixel losses, such as the mean squared error between predicted and ground truth density maps. In the few-shot (or reference-based) counting setting, each image $I_{i}$ is paired with $K$ exemplars, defined as bounding boxes $B^{E}={b_{i}}_{i=1:K}\in\mathbb{R}^{4}$ that depict instances of the target class within the same image [32], with $K=3$ in the original formulation. Object classes are split into base $C_{b}$ and novel $C_{n}$ categories, with $C_{b}\cap C_{n}=\emptyset$. For $C_{b}$, both density maps and exemplars are available; for $C_{n}$, only exemplars are provided, and the goal is to count instances by transferring knowledge from $C_{b}$ (see also Fig. 4). In contrast, reference-less and text-guided zero-shot settings omit exemplars: the former automatically infers dominant class(es) via instance repetition, while the latter uses a textual description $T_{i}$ to specify the target class for each image $I_{i}$.